The solution of D.E(D^4-16)y=0 is
1
Expert's answer
2020-12-23T17:56:36-0500

"y^{(iv)} - 16y = 0\\
\textsf{The auxiliary equation is}\,\,\,m^4 - 16 = 0.\\

(m^2 - 4)(m^2 + 4) = 0\\

m^2 = 4, m^2 = -4\\

m = \pm 2, \pm 2i.\\

\textsf{Recall that if the solution to the}\\
\textsf{auxiliary equation is of the form}\,\,\, \alpha + i\beta, \\
\textsf{the solution is of the form} \\
y = e^{\alpha x}\left(A\cos(\beta x) + B\sin(\beta x)\right).\\

\therefore y = Ae^{2x} + Be^{-2x} + C\cos(2x) + D\sin(2x).\,\,\,\textsf{is a solution to the ODE}"